Is there a solution for:
23-t <= 2 (t-9) -3 (t+2)
thanks, Donnie

Answer by stanbon(48546)

Is there a solution for:
23-t <= 2 (t-9) -3 (t+2)
23 - t <= 2t - 18 -3t - 6
23 - t <= -t - 24
23 < = -24
That is false, so there is no solution fot the original equation
=============================
Cheers,
Stan H.